OpenStudy (anonymous):

Which equation represents the relationship shown in the table below? (1 point) y = –x – 3 y = x – 3 y = 2x − 3 y = ­­–2x + 3

OpenStudy (anonymous):

theres the table

OpenStudy (anonymous):

can you place another table , cant be opened!

OpenStudy (anonymous):

is that better ?

OpenStudy (anonymous):

yeah wait

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

y=2x-3

OpenStudy (anonymous):

Thanks a lot :) im trying to get caught up i started school late

OpenStudy (anonymous):

formula : y-y0=slope(x-x0) slope=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw